A piece of iron weighs 100gm in air, 90gm in water.what is the specific gravity of water?
1
Expert's answer
2012-06-22T08:56:18-0400
delta(P) = pw * V * g
V = m / pi = 0.1 / 7874
pw = delta(P) / (V *
g)
pw = (0.100 - 0.090) / ((0.1 / 7874) * 9.81) = 80.265